3r-4s=0 and 2r+5s=23
Answer by checkley75(3666)

3R-4S=0
3R=4S
R=4S/3 NOW SUBSTITUTE (4S/3) FOR R ANS SOLVE FOR S
2(4S/3)+5S=23
12S/3+5S=23
(12S+3*5S)/3=23 NOW CROSS MULTIPLY
12S+15S=3*23
27S=69
S=69/27
S=23/9 ANSWER.
R=(4*23/9)/3
R=92/9*1/3
R=92/27